A curve is such that dy/dx = 12/ (2x+1)squared . The point (1 , 1) lies on the curve .find the coordinates of the point at which the curve intersects the x axis
1
Expert's answer
2019-09-02T13:50:44-0400

"\frac{}{}" "\frac{dy}{dx }= \frac{12}{ (2x+1)^2}"


"dy= \frac{12{dx }}{ (2x+1)^2}"

"y= \int\frac{12{dx }}{ (2x+1)^2}=12\int\frac{{dx }}{ (2x+1)^2}=6\int\frac{{d (2x) }}{ (2x+1)^2}=\\= 6\int\frac{{d (2x+1) }}{ (2x+1)^2}"


"t=2x+1"

"\\ 6\int\frac{{d (2x+1) }}{ (2x+1)^2}= 6\int\frac{{d t }}{ t^2}= \frac{{-6 }}{ t}+C= \frac{-6 }{(2x+1) }+C"


"y= - \frac{6 }{(2x+1) }+C"


The point (1 , 1) lies on the curve:

"y(1)=1 \Rightarrow -\frac{6 }{(2*1+1) }+C=1 \Rightarrow \\C=1 -\frac{-6 }{(2*1+1) } \Rightarrow \\ 
\Rightarrow C=1 +2 \Rightarrow C=3"


"y= \frac{-6 }{2x+1 }+3 \\"


"y= \frac{6x-3 }{2x+1 }"

The coordinates of the point at which the curve intersects the x axis


"\frac{6x-3 }{2x+1 }=0 \Rightarrow {6x-3=0 } ( 2x+1 \neq 0 ) \Rightarrow x=\frac{1 }{2 }" .


Answer 

The coordinates of the point at which the curve "y= \frac{6x-3 }{2x+1 }" intersects the x axis:

"x=\frac{1 }{2 }" .